OpenStudy (alli6106):

Eric's gas tank is 12 full. After he buys 3 gallons of gas, it is 23 full. How many gallons can Eric's tank hold?

OpenStudy (danjs):

Let T = # gallons in a full tank Halft a full tank + 3 gallons = 2/3 a full tank \[\frac{ 1 }{ 2 }T + 3 = \frac{ 2 }{ 3 }T\]

OpenStudy (danjs):

multiply everything by 6 first

OpenStudy (danjs):

I am assuming those are 1/2 and 2/3 of a tank in the problem, you forgot the division symbol

OpenStudy (danjs):

The answer is not 40 gallons like above.

OpenStudy (alli6106):

WHAT IS THE ANSWER